# TOP2024: an overview of experimental results

### v1

@@ 1 Introduction
Since its discovery in 1995 [1, 2], the heaviest known particle, the top quark, has been a focus of studies in the experimental and theoretical particle physics. The large mass of top quark results in a very short life time as well as a Yukawa coupling close to one. The particle is therefore an excellent candidate to study the bare quark properties, to understand the electroweak symmetry breaking, and to connect with new phenomena. In proton-proton (pp) collisions at the CERN LHC, top quarks are abundantly produced in pair ($\text{t}\bar{\text{t}}$) through strong interactions, and to lesser extent, single via electroweak processes (single-top). The production in both cases can be accompanied by extra particles, including jets, vector bosons, the Higgs boson, and even additional top quarks. Currently, most of the production modes are observed and/or precisely measured. From precise measurements, the top quark mass and properties are determined while additional parameters of the standard model of particle physics (SM) are extracted. At the same time, searches for new phenomena with top quarks together with the interpretation of top quark measurements have put strong constraints on theories beyond the standard model (BSM).
@@
Considering different top quark production modes, the large cross section of $\text{t}\bar{\text{t}}$, $\sigma_{\text{t}\bar{\text{t}}}$, has made it a unique target for the additional measurements and interpretations. It was the first to measure at the start of the LHC Run 3 [5, 6], already with a high precision. In addition, CMS reported a measurements at $\sqrt{s}=5.02\,\text{TeV}$ [7], to compare with the corresponding ATLAS result [8] and to be used as a reference for top quark studies in heavy ion collisions at the same center-of-mass energy, $E_{\text{c.o.m}}$. The ATLAS collaboration presented a $\sigma_{\text{t}\bar{\text{t}}}$ measurements in proton-lead (pPb) collisions at $\sqrt{s}=8.16\,\text{TeV}$ [9] to compare with an earlier result by CMS [10]. New results from both experiments call for further discussions and comparisons of analysis details and simulation settings, which can be particularly interesting when heavy ions are involved.
The available knowledge and expertise paves the way for future investigations in the field of top quark physics. By the end of 2025, the full LHC data at $\sqrt{s}=13.6\,\text{TeV}$ will be available for physics analyses. The High-Luminosity LHC is expected to start in , providing the largest-ever amount of data. At the same time, experiments will undergo a substantial upgrade to sustain the immense luminosity and to fully exploit the wealth of the data. This is a unique opportunity to explore uncharted territories in top quark physics, drawing on the existing and developing knowledge.
@@ 3 Opportunities with $\text{t}\bar{\text{t}}$ differential measurements
Differential distributions of $\sigma_{\text{t}\bar{\text{t}}}$ have been measured since the early times of the LHC data taking and provided valuable inputs to the upcoming measurements and to the theory community. The well-known observed slope in the differential measurement of top quark transverse momentum $p_{\text{T}}$ with respect to theory predictions at the time triggered a chain of studies over different data-taking eras. Multidifferential measurements were performed to possibly localize the effect where at the same time, higher order calculations could explain (part of) the disagreements. The acquired knowledge from multidifferential measurements went beyond understanding the top quark $p_{\text{T}}$ distributions and used to extract SM parameters such as $\alpha_{\text{S}}$ and top quark mass, and to constrain the gluon density function of protons  [^3] Refer to this presentation for a comprehensive overview.. The state-of-the-art ML methods have made it possible to perform simultaneous differential measurements of many more observables with a higher granularity in other physics processes [16, 17]. It is desired to use the same techniques for differential $\sigma_{\text{t}\bar{\text{t}}}$ measurements.
@@
Heavy pseudoscalars decaying to $\text{t}\bar{\text{t}}$ can create a peak-dip structure in the $m_{\text{t}\bar{\text{t}}}$ distributions, if exist. In a search for such particles, using $\text{t}\bar{\text{t}}$ angular and spin correlation observables, CMS has observed an excess with a significance of more than $5\sigma$ at $m_{\text{t}\bar{\text{t}}}$ values close to the $\text{t}\bar{\text{t}}$ production threshold [18]. The excess is compatible with the presence of a toponium signal, using the available calculations in which the color-octet state is not accounted for. Figure 2 shows the observed and expected $m_{\text{t}\bar{\text{t}}}$ distributions in bins of spin correlation observable. The SM distributions with an injected topnium signal are fitted to and compared with the data in the lower panel. The pseudoscalar hypothesis is slightly favored by the data for the observed excess. A similar measurement by ATLAS together with more realistic theory predictions and proposals for complementary measurements seem necessary to better identify the nature of this signal.
@@
Using the $\Delta\phi$ between the two leptons from W bosons in $\text{t}\bar{\text{t}}$ decays, ATLAS and CMS experiments observed quantum entanglement between the two top quarks near the $\text{t}\bar{\text{t}}$ production threshold [19, 20]. The entanglement signal remained statistically significant in the presence of possible $\text{t}\bar{\text{t}}$ bound state. In another work [21], the CMS experiment exploited the full $\text{t}\bar{\text{t}}$ spin density matrix and observed entangled top quarks that in part of the phase space with high $m_{\text{t}\bar{\text{t}}}$ and low top quark scattering angles. This is particularly important because in this regions top quarks are expected to be largely outside each other’s light cone and the information cannot be transmitted with hidden variables  [^4] Possibilities to experimentally probe quantum entanglement in single-top were presented here..
@@ 4 Associated top quark productions
While the latest $\text{t}\bar{\text{t}}$$\text{b}\bar{\text{b}}$ measurements by ATLAS and CMS [22, 23] show similar trends in comparison with theory predictions, comparing of the results is not straight forward because of different final states and definitions of the fiducial phase space (see Fig. 3 for the ATLAS result). In addition to harmonized selections, common MC samples with similar generators and scale choices would be extremely useful in this context. The CMS $\text{t}\bar{\text{t}}$$\text{c}\bar{\text{c}}$ measurement has to be updated with the high-performance c-taggers, and in view of the latest ATLAS result [24]. The LHC searches for $\text{t}\bar{\text{t}}$$\text{t}\bar{\text{t}}$ have been concluded with the observation of the process in Run 2, with the possibility to combination the results of the two experiments with higher precisions in Run 3. The off-shell $\text{t}\bar{\text{t}}$$\text{t}\bar{\text{t}}$ production has shown constraining power on the Higgs boson width and top-Higgs Yukawa coupling $\kappa_{\text{t}}$ [25]. Given the sensitivity of differential $m_{\text{t}\bar{\text{t}}}$ and $\Delta y_{\text{t}\bar{\text{t}}}$ measurements to $\kappa_{\text{t}}$ [26], there is more room for top quark physics to contribute to the Higgs sector of the SM.
Figure 3: Measured fiducial cross sections for compared with the central values of NLO $\text{t}\bar{\text{t}}$$\text{b}\bar{\text{b}}$ predictions from different event generators and settings [22]. The right hand panel shows the ratios of the MC predictions to the data. A similar trend is observed in the CMS measurement [23] performed in lepton+jets final states and compared with different event generators and settings.
@@ 6 Conclusion
The top quark physics at the LHC advances relying on the outstanding work from theory and experiment. Exchanges across the communities are vital for the continuation of the research. On the experimental side, the succesful experience of common Monte-Carlo $\text{t}\bar{\text{t}}$ samples between ATLAS and CMS can be extended to other top quark processes to facilitate the comparison of definitions and the results.

### v2

@@ 1 Introduction
Since its discovery in 1995 [1, 2], the heaviest known particle, the top quark, has been a focus of studies in the experimental and theoretical particle physics. The large mass of the top quark results in a very short life time as well as a Yukawa coupling close to one. The particle is therefore an excellent candidate to study the bare quark properties, to understand the electroweak symmetry breaking, and to connect with new phenomena. In proton-proton (pp) collisions at the Large Hadron Collider (LHC), top quarks are abundantly produced in pair ($\text{t}\bar{\text{t}}$) through strong interactions, and to lesser extent, single via electroweak processes (single-top). The production in both cases can be accompanied by extra particles, including jets, vector bosons, the Higgs boson, and even additional top quarks. Currently, most of the production modes are observed and/or precisely measured. From precise measurements, the top quark mass and properties are determined while some other parameters of the standard model of particle physics (SM) such as the strong coupling constant, $\alpha_{\text{S}}$, can be extracted. At the same time, searches for new phenomena with top quarks together with the interpretation of top quark measurements have put strong constraints on theories beyond the standard model (BSM).
@@
Considering different top quark production modes, the large $\text{t}\bar{\text{t}}$ cross section, $\sigma_{\text{t}\bar{\text{t}}}$, has made it a unique target for the measurement of top quark properties, extraction of relevant SM parameters, and interpretations in the context of new theories. It was the first to be measured at the start of the LHC Run 3 [5, 6], already with a high precision. In addition, CMS reported a measurement at $\sqrt{s}=5.02\,\text{TeV}$ [7], to compare with the corresponding ATLAS result [8] and to be used as a reference for top quark studies in heavy ion collisions at the same center-of-mass energy, $E_{\text{c.o.m}}$. The ATLAS collaboration presented a $\sigma_{\text{t}\bar{\text{t}}}$ measurements in proton-lead (pPb) collisions at $\sqrt{s}=8.16\,\text{TeV}$ [9] to compare with an earlier result by CMS [10]. New results from both experiments call for further discussions and comparisons of analysis details and simulation settings, which can be particularly interesting when heavy ions are involved.
The available knowledge and expertise paves the way for future investigations in the field of top quark physics. By the end of 2025, the full LHC data at $\sqrt{s}=13.6\,\text{TeV}$ will be available for physics analyses. The High-Luminosity LHC is expected to start in mid-2030, providing the largest-ever amount of proton-proton collision data. At the same time, experiments will undergo a substantial upgrade to sustain the immense luminosity and to fully exploit the wealth of the data. This is a unique opportunity to explore uncharted territories in top quark physics, drawing on the existing and developing knowledge.
@@ 3 Opportunities with $\text{t}\bar{\text{t}}$ differential measurements
Differential distributions of $\sigma_{\text{t}\bar{\text{t}}}$ have been measured since the early times of the LHC data taking and provided valuable inputs to the upcoming measurements and to the theory community. The well-known observed slope in the differential measurement of top quark transverse momentum $p_{\text{T}}$ with respect to theory predictions at the time triggered a chain of studies over different data-taking eras. Multidifferential measurements were performed to possibly localize the effect where at the same time, higher order calculations could explain (part of) the disagreements. The acquired knowledge from multidifferential measurements went beyond understanding the top quark $p_{\text{T}}$ distributions and are used to extract SM parameters such as $\alpha_{\text{S}}$ and top quark mass, and to constrain the gluon density function of protons  [^3] Refer to this presentation for a comprehensive overview.. The state-of-the-art ML methods have made it possible to perform simultaneous differential measurements of many more observables with a higher granularity in other physics processes [16, 17]. It is desired to use the same techniques for differential $\sigma_{\text{t}\bar{\text{t}}}$ measurements.
@@
Heavy pseudoscalars decaying to $\text{t}\bar{\text{t}}$ can create a peak-dip structure in the $m_{\text{t}\bar{\text{t}}}$ distributions, if exist. In a search for such particles, using $\text{t}\bar{\text{t}}$ angular and spin correlation observables, CMS has observed an excess with a significance of more than $5\sigma$ at $m_{\text{t}\bar{\text{t}}}$ values close to the $\text{t}\bar{\text{t}}$ production threshold [18]. The excess is compatible with the presence of a toponium signal, using the available calculations in which the color-octet state is not accounted for. Figure 2 shows the observed and expected $m_{\text{t}\bar{\text{t}}}$ distributions in bins of spin correlation observable. The SM distributions with an injected toponium signal are fitted to and compared with the data in the lower panel. The pseudoscalar hypothesis is slightly favored by the data for the observed excess. A similar measurement by ATLAS together with more realistic theory predictions and proposals for complementary measurements seem necessary to better identify the nature of this signal.
@@
Using the $\Delta\phi$ between the two leptons from W bosons in $\text{t}\bar{\text{t}}$ decays, ATLAS and CMS experiments observed quantum entanglement between the two top quarks near the $\text{t}\bar{\text{t}}$ production threshold [19, 20]. The entanglement signal remained statistically significant in the presence of possible $\text{t}\bar{\text{t}}$ bound state. In another work [21], the CMS experiment exploited the full $\text{t}\bar{\text{t}}$ spin density matrix and observed entangled top quarks in part of the phase space with high $m_{\text{t}\bar{\text{t}}}$ and low top quark scattering angles. This is particularly important because in this regions top quarks are expected to be largely outside each other’s light cone and the information cannot be transmitted with hidden variables  [^4] Possibilities to experimentally probe quantum entanglement in single-top were presented here..
@@ 4 Associated top quark productions
While the latest $\text{t}\bar{\text{t}}$$\text{b}\bar{\text{b}}$ measurements by ATLAS and CMS [22, 23] show similar trends in comparison with theory predictions, comparing of the results is not straight forward because of different final states and definitions of the fiducial phase space (see Fig. 3 for the ATLAS result). In addition to harmonized selections, common MC samples with similar generators and scale choices would be extremely useful in this context. The CMS $\text{t}\bar{\text{t}}$$\text{c}\bar{\text{c}}$ measurement has to be updated with the high-performance c-taggers, and in view of the latest ATLAS result [24]. The LHC searches for $\text{t}\bar{\text{t}}$$\text{t}\bar{\text{t}}$ have been concluded with the observation of the process in Run 2, with the possibility to combine the results of the two experiments with higher precisions in Run 3. The off-shell $\text{t}\bar{\text{t}}$$\text{t}\bar{\text{t}}$ production has shown constraining power on the Higgs boson width and top-Higgs Yukawa coupling $\kappa_{\text{t}}$ [25]. Given the sensitivity of differential $m_{\text{t}\bar{\text{t}}}$ and $\Delta y_{\text{t}\bar{\text{t}}}$ measurements to $\kappa_{\text{t}}$ [26], there is more room for top quark physics to contribute to the Higgs sector of the SM.
Figure 3: Measured fiducial cross sections compared with the central values of NLO $\text{t}\bar{\text{t}}$$\text{b}\bar{\text{b}}$ predictions from different event generators and settings [22]. The right hand panel shows the ratios of the MC predictions to the data. A similar trend is observed in the CMS measurement [23] performed in lepton+jets final states and compared with different event generators and settings.
@@ 6 Conclusion
Top quark physics at the LHC advances relying on the outstanding work from theory and experiments. Exchanges across the communities are vital for the continuation of the research. On the experimental side, the succesful experience of common Monte-Carlo $\text{t}\bar{\text{t}}$ samples between ATLAS and CMS can be extended to other top quark processes to facilitate the comparison of definitions and the results.